Solve the system by elimination.
-2x+2y+3z=0
-2x-y+z=-3
2x+3y+3z=5

Answer by Alan3354(69443)

-2x+2y+3z=0
-2x-y+z=-3
2x+3y+3z=5
---------------
-2x+2y+3z=0
-2x-y+z=-3
---------------- Subtract
3y + 2z = 3 Eqn A
=========================
-2x-y+z=-3
2x+3y+3z=5
---------------- Add
2y + 4z = 2
6y + 4z = 6 Eqn A times 2
------------------------------------ Subtract
-4y = -4
y = 1
-----------------
z = 0
-------------------
x = 1